OpenStudy (anonymous):

I think this problem will require later to solve a fourth degree equation. Is there any shorter method for this?

OpenStudy (anonymous):

If \[\sqrt{m}+n=3\]and \[m+\sqrt{n}=5\] find the values of m and n.

terenzreignz (terenzreignz):

Maybe, maybe not... have you done using using 'conventional' methods?

terenzreignz (terenzreignz):

Just one 'using'... gah

terenzreignz (terenzreignz):

<shrugs> let's see... \[\Large n = 3- \sqrt m\]\[\Large m + \sqrt{3-\sqrt{m}}= 5\]\[\Large \sqrt{3-\sqrt m}= 5- m\]Squaring both sides\[\Large3-\sqrt m = m^2 -10m + 25\]\[\Large m^2 -10m + \sqrt m +22= 0\]

terenzreignz (terenzreignz):

This looks highly unappealing... \[\Large u = \sqrt m\]\[\Large u^4 -10u^2 + u + 22 = 0\]

terenzreignz (terenzreignz):

Come on, bro, make your presence felt :D